# MIT-CTP/5902

### v1

@@ C School of Natural Sciences, Institute for Advanced Study, Princeton, NJ
LSM and CPT
Nathan SeibergC, Shu-Heng ShaoP, and Wucheng ZhangT
@@ 1.4 Back to the continuum: emanant symmetries and emergent anomalies
This emanant $\mathbb{Z}_{n}$ symmetry is not a subgroup of the exact $\mathbb{Z}$ (or $\mathbb{Z}_{L}$) lattice translation symmetry. Instead, it is a quotient of it. Therefore, the IR theory can be placed in a background $\mathbb{Z}_{n}$ gauge field that does not correspond to any background of the underlying UV lattice theory. In this background, the IR theory can have anomalies that do not correspond to any UV anomaly. Such anomalies are referred to as emergent anomalies [23]. See also [39, 40, 41].
@@ 2.1 Lattice translation and CRT
In this section, we assume $N_{f}L$ to be even, so that the total number of fermions is even. [^6] There are various subtleties associated with having an odd number of fermions. See e.g., [50, 51, 52, 29, 53] for recent discussions. In particular, for odd $N_{f}L$, the fermion parity is an outer automorphism, and there is no unitary operator on the Hilbert space implementing this symmetry transformation. As a result, the Hilbert space is not graded. Also, the standard relation between this Hilbert space and the path integral is not maintained. In our case, when $N_{f}=3,7$ mod 8 and $L$ is odd, the anti-unitary symmetry $\Pi$ defined below is also an outer automorphism. See Appendices B and D.2. In this case, we can define the fermion parity operator:
@@
$\Pi\,\chi_{\ell}^{A}\,\Pi^{-1}=\chi_{-\ell+1}^{A}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Pi\,i\,\Pi^{-1}=-i\,.$ (2.8)
@@
[…]
@@
[…]
@@ 2.2 Lattice Hamiltonians and the continuum limit
$\Theta\,\chi_{\text{L}}^{A}(t,x)\,\Theta^{-1}=-\chi_{\text{L}}^{A}(-t,-x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Theta\,\chi_{\text{R}}^{A}(t,x)\,\Theta^{-1}=\chi_{\text{R}}^{A}(-t,-x)\,.$ (2.18)
@@ 2.3.1 Anomaly for ${\tilde{\Pi}}=\mathcal{R}\mathcal{T}$
${\tilde{\Pi}}\left(\chi^{A}_{\ell}+\chi^{A}_{-\ell}\right){\tilde{\Pi}}^{-1}=\left(\chi^{A}_{\ell}+\chi^{A}_{-\ell}\right)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\tilde{\Pi}}\left(\chi^{A}_{\ell}-\chi^{A}_{-\ell}\right){\tilde{\Pi}}^{-1}=-\left(\chi^{A}_{\ell}-\chi^{A}_{-\ell}\right)\,$, (2.21)
@@ 2.3.2 Anomaly between translation and $\Pi$ (the lattice CRT)
$\Pi\,T=(-1)^{\frac{N_{f}(N_{f}-1)}{2}}T^{-1}\,\Pi\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{even}\leavevmode\nobreak\ L\,.$ (2.28)
@@
$\Pi^{2}=(-1)^{\frac{N_{f}(N_{f}-2)}{8}},\quad\Pi(-1)^{F}=(-1)^{\frac{N_{f}}{2}}(-1)^{F}\Pi\,,\quad\text{odd}\leavevmode\nobreak\ L\,.$ (2.29)
@@ 2.3.3 Matching the lattice anomalies with the continuum
We now compare these lattice anomalies of the gapless Majorana chain (2.15) with their continuum counterparts. In the continuum field theory in (2.16), the fermion parity $(-1)^{F}$ and the chiral fermion parity $\mathsf{C}$, which emanates from lattice translation $T$, have an order 8 anomaly classified by $\text{Hom(}\,\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2}),U(1))=\mathbb{Z}_{8}$[64]. (See [65, 66] for earlier works on 2+1d fermionic SPT phases with a unitary $\mathbb{Z}_{2}$ symmetry.) On the lattice, however, we only find an order 2 anomaly between $T$ and $(-1)^{F}$. This difference arises because the symmetry algebras are not the same: the continuum $\mathbb{Z}_{2}$ symmetry is a quotient of the lattice translation group $\mathbb{Z}_{L}$ for even $L$. Consequently, the continuum $\mathbb{Z}_{2}$ symmetry is an emanant symmetry and its anomaly is an emergent anomaly. (See Section 1.4.) This leads to a natural question. Can we detect this mod 8 emergent anomaly on the lattice? In fact, it was argued in [67] that this mod 8 anomaly cannot be realized exactly on the Majorana chain by a unitary, locality-preserving, $\mathbb{Z}_{2}$ lattice operator. (See the same reference for a realization of such a lattice operator on a non-tensor product Hilbert space.)
@@ 2.4 Lattice realization of the Smith isomorphism
For all $L$, these zero modes are robust and are protected by $\tilde{\Pi}$}: they cannot be lifted by local, $\tilde{\Pi}$}-symmetric deformation of the Hamiltonian unless $N_{f}=0$ mod 8.131313One might suspect that the protected Majorana zero modes arise solely because the two sides of the interfaces belong to different invertible phases of [72], and are thus unrelated to $\tilde{\Pi}$}. This, however, is not the case. Imposing $\tilde{\Pi}$} leads to a stronger conclusion. First, the two sides correspond to distinct invertible phases only when $N_{f}$ is odd. Second, even in this case, the phase difference protects the number of zero modes only modulo 2, whereas $\tilde{\Pi}$} protects this number modulo 8. When $N_{f}=0$ mod 8, we can add an onsite, $\tilde{\Pi}$}-symmetric quartic term $\sum_{A,B,C,D}\mu_{ABCD}\chi^{A}_{\ell}\chi^{B}_{\ell}\chi^{C}_{\ell}\chi^{D}_{\ell}$ at the interface to lift the zero modes. This is consistent with the discussion of the modulo 8 anomaly for $\tilde{\Pi}$} in Section 2.3.1.
@@ 3.1 Lattice translation and CRT
$T\,\vec{S}_{j}\,T^{-1}=\vec{S}_{j+1}\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{R}\,\vec{S}_{j}\,\mathcal{R}^{-1}=\vec{S}_{-j}\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{T}\,\vec{S}_{j}\,\mathcal{T}^{-1}=-\vec{S}_{j}\,.$ (3.1)
Here $T$ and $\mathcal{R}$ are unitary operators, while the time-reversal operator $\mathcal{T}$ is anti-unitary, i.e., $\mathcal{T}i\mathcal{T}^{-1}=-i$. Explicitly, $\mathcal{T}=\left(\prod_{j=1}^{L}iY_{j}\right)\mathcal{K}$, where $\mathcal{K}$ is the complex conjugation. [^14] The anti-unitarity of $\mathcal{T}$ forces the minus sign in its action on $\vec{S}_{j}$, such that the commutation relations of $\vec{S}_{j}$ are preserved. Related to that, $\mathcal{T}$ acts on each qubit such that its square is 1-1, i.e., each qubit is a Kramers’ doublet. On the other hand, $T$ and $\mathcal{R}$ are products of the SWAP gates. These operators and their LSM constraints were discussed in [73, 74, 75, 76, 63]. Below, we will discuss a related, but different anomaly associated with lattice translation $T$ and the following anti-unitary operator
@@
$T^{L}=1\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Pi^{2}=(-1)^{L}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ T\Pi=\Pi T^{-1}\,.$ (3.4)
The important $(-1)^{L}$ phase follows from the fact that each qubit is a Kramers’ doublet (see footnote 14). For completeness, we also record the larger algebra generated by $\mathcal{R}$, $\mathcal{T}$, and $T$:
@@ 3.2.1 Anti-ferromagnetic phase
${\cal L}={f\over 2}\partial_{\mu}\vec{n}\cdot\partial^{\mu}\vec{n}+\theta\Gamma\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \vec{n}\cdot\vec{n}=1$ (3.11)
@@
$\mathsf{C}\,\vec{n}(t,x)\,\mathsf{C}^{-1}=-\vec{n}(t,x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{R}\,\vec{n}(t,x)\,\mathsf{R}^{-1}=\vec{n}(t,-x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{T}\,\vec{n}(t,x)\,\mathsf{T}^{-1}=-\vec{n}(-t,x)\,.$ (3.12)
@@
In the quantum theory, the classical field $\vec{n}$ becomes an operator, which is not constrained by $\vec{n}^{2}=1$. (In order not to clutter the equations, we will not distinguish between the classical field and its quantum counterpart.) For $\theta\in\pi(1+2{\mathbb{Z}})$, this continuum theory flows to the $su(2)_{1}$ CFT, which is the same as the IR limit of the lattice model (3.9). The lattice operators $\vec{S}_{j}$ flow to the continuum operators [77]
@@ 3.2.2 Ferromagnetic phase
$\mathsf{R}\,\vec{n}(t,x)\,\mathsf{R}^{-1}=\vec{n}(t,-x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{T}\,\vec{n}(t,x)\,\mathsf{T}^{-1}=-\vec{n}(-t,x)$ (3.17)
@@ 3.3.1 Anomaly for ${\tilde{\Pi}}=\mathcal{R}\mathcal{T}$
${\tilde{\Pi}}^{2}=-1\,,\quad\text{odd}\leavevmode\nobreak\ L\,.$ (3.19)
@@ 3.3.2 Anomaly between translation and $\Pi$ (the lattice CRT)
$\Pi^{2}=-1\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{odd }L\,$, (3.22)
@@ 3.4 Lattice realization of the Smith isomorphism
As discussed in Appendix F.4, this deformation gaps the bulk of the two halves of the circle. Indeed, under the identification $(-1)^{j}\vec{S}_{j}\cdot\vec{S}_{j+1}\to\Gamma$ [77] (see Appendix F.2), (LABEL:HDWs) corresponds to changing the $\theta$-angle by the opposite amounts away from $\theta=\pi$ on the left and right halves of the circle.
@@ 4.1 The model and its symmetries
We are particularly interested in a $\mathbb{Z}_{2}^{\mathcal{U}}$ subgroup of the internal symmetry, which is generated by [^20] When $L=0$ mod 4, one can perform a unitary transformation $\prod_{j=1,2\leavevmode\nobreak\ \text{mod}\leavevmode\nobreak\ 4}X_{j}$ to map $\mathcal{U}$ to $\prod_{j=1}^{L}\mathrm{CZ}_{j,j+1}\prod_{j=1}^{L}X_{j}$, which is sometimes known as the CZX symmetry [92].

### v2

@@ C School of Natural Sciences, Institute for Advanced Study, Princeton, NJ
We study a number of 1+1d lattice models with anti-unitary symmetries that simultaneously reflect space and reverse time. Some of these symmetries are anomalous, leading to Lieb-Schultz-Mattis-type constraints, thus excluding a trivially gapped phase. Examples include a mod 8 anomaly in the Majorana chain and various mod 2 anomalies in the spin chain. In some cases, there is an exact, non-anomalous lattice symmetry that flows in the continuum to CPT. In some other cases, the CPT symmetry of the continuum theory is emergent or absent. Depending on the model, the anomaly of the lattice model is matched in the continuum in different ways. In particular, it can be mapped to an emergent anomaly of an emanant symmetry.
@@ 1.4 Back to the continuum: emanant symmetries and emergent anomalies
This emanant $\mathbb{Z}_{n}$ symmetry is not a subgroup of the exact $\mathbb{Z}$ (or $\mathbb{Z}_{L}$) lattice translation symmetry. Instead, it is a quotient of it. Therefore, the IR theory can be placed in a background $\mathbb{Z}_{n}$ gauge field that does not correspond to any background of the underlying UV lattice theory.
Thus, the emanant symmetry in the IR may have a nontrivial ’t Hooft anomaly even when the corresponding UV symmetry is anomaly-free. In such cases, although the symmetry itself is emanant, its anomaly emerges out of nowhere. Such anomalies are therefore referred to as emergent anomalies [23]. See also [39, 40, 41].
@@ 2.1 Lattice translation and CRT
In this section, we assume $N_{f}L$ to be even, so that the total number of fermions is even. [^6] There are various subtleties associated with having an odd number of fermions. See e.g., [50, 51, 52, 29, 53, 54] for recent discussions. In particular, for odd $N_{f}L$, the fermion parity is an outer automorphism, and there is no unitary operator on the Hilbert space implementing this symmetry transformation. As a result, the Hilbert space is not graded. Also, the standard relation between this Hilbert space and the path integral is not maintained. In our case, when $N_{f}=3,7$ mod 8 and $L$ is odd, the anti-unitary symmetry $\Pi$ defined below is also an outer automorphism. See Appendices B and D.2. In this case, we can define the fermion parity operator:
@@
$\Pi\,\chi_{\ell}^{A}\,\Pi^{-1}=\chi_{-\ell+1}^{A}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Pi\,i\,\Pi^{-1}=-i\,.$ (2.8)
@@
$[(-1)^{F}]^{2}=1\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$T^{L}=1\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$\Pi^{2}=1\,$, (2.9)
$(-1)^{F}\Pi=\Pi(-1)^{F}\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$T(-1)^{F}=(-1)^{N_{f}}(-1)^{F}T\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$\Pi\,T=(-1)^{\frac{N_{f}(N_{f}-1)}{2}}T^{-1}\,\Pi\,.$
@@
$[(-1)^{F}]^{2}=1\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$T^{L}=1\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$\Pi^{2}=(-1)^{\frac{N_{f}(N_{f}-2)}{8}},$ (2.10)
@@ 2.2 Lattice Hamiltonians and the continuum limit
$\Theta\,\chi_{\text{L}}^{A}(t,x)\,\Theta^{-1}=-\chi_{\text{L}}^{A}(-t,-x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Theta\,\chi_{\text{R}}^{A}(t,x)\,\Theta^{-1}=\chi_{\text{R}}^{A}(-t,-x)\,.$ (2.18)
@@ 2.3.1 Anomaly for ${\tilde{\Pi}}=\mathcal{R}\mathcal{T}$
${\tilde{\Pi}}\left(\chi^{A}_{\ell}+\chi^{A}_{-\ell}\right){\tilde{\Pi}}^{-1}=\left(\chi^{A}_{\ell}+\chi^{A}_{-\ell}\right)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\tilde{\Pi}}\left(\chi^{A}_{\ell}-\chi^{A}_{-\ell}\right){\tilde{\Pi}}^{-1}=-\left(\chi^{A}_{\ell}-\chi^{A}_{-\ell}\right)\,$, (2.21)
@@ 2.3.2 Anomaly between translation and $\Pi$ (the lattice CRT)
$\Pi\,T=(-1)^{\frac{N_{f}(N_{f}-1)}{2}}T^{-1}\,\Pi\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{even}\penalty 10000\ L\,.$ (2.28)
@@
$\Pi^{2}=(-1)^{\frac{N_{f}(N_{f}-2)}{8}},\quad\Pi(-1)^{F}=(-1)^{\frac{N_{f}}{2}}(-1)^{F}\Pi\,,\quad\text{odd}\penalty 10000\ L\,.$ (2.29)
@@ 2.3.3 Matching the lattice anomalies with the continuum
We now compare these lattice anomalies of the gapless Majorana chain (2.15) with their continuum counterparts. In the continuum field theory in (2.16), the fermion parity $(-1)^{F}$ and the chiral fermion parity $\mathsf{C}$, which emanates from lattice translation $T$, have an order 8 anomaly classified by $\text{Hom(}\,\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2}),U(1))=\mathbb{Z}_{8}$[65, 66, 67]. [^11] See [68] for earlier work on 2+1d fermionic SPT phases with a unitary $\mathbb{Z}_{2}$ symmetry. See also [69, 70] for another approach to fermionic SPT phases using operator algebras on infinite lattices and the importance of the lattice CRT symmetry. On the lattice, however, we only find an order 2 anomaly between $T$ and $(-1)^{F}$. This difference arises because the symmetry algebras are not the same: the continuum $\mathbb{Z}_{2}$ symmetry is a quotient of the lattice translation group $\mathbb{Z}_{L}$ for even $L$. Consequently, the continuum $\mathbb{Z}_{2}$ symmetry is an emanant symmetry and its anomaly is an emergent anomaly. (See Section 1.4.) This leads to a natural question. Can we detect this mod 8 emergent anomaly on the lattice? In fact, it was argued in [71] that this mod 8 anomaly cannot be realized exactly on the Majorana chain by a unitary, locality-preserving, $\mathbb{Z}_{2}$ lattice operator. (See the same reference for a realization of such a lattice operator on a non-tensor product Hilbert space.)
@@ 2.4 Lattice realization of the Smith isomorphism
For all $L$, these zero modes are robust and are protected by $\tilde{\Pi}$}: they cannot be lifted by local, $\tilde{\Pi}$}-symmetric deformation of the Hamiltonian unless $N_{f}=0$ mod 8.141414One might suspect that the protected Majorana zero modes arise solely because the two sides of the interfaces belong to different invertible phases of [76], and are thus unrelated to $\tilde{\Pi}$}. This, however, is not the case. Imposing $\tilde{\Pi}$} leads to a stronger conclusion. First, the two sides correspond to distinct invertible phases only when $N_{f}$ is odd. Second, even in this case, the phase difference protects the number of zero modes only modulo 2, whereas $\tilde{\Pi}$} protects this number modulo 8. When $N_{f}=0$ mod 8, we can add an onsite, $\tilde{\Pi}$}-symmetric quartic term $\sum_{A,B,C,D}\mu_{ABCD}\chi^{A}_{\ell}\chi^{B}_{\ell}\chi^{C}_{\ell}\chi^{D}_{\ell}$ at the interface to lift the zero modes. This is consistent with the discussion of the modulo 8 anomaly for $\tilde{\Pi}$} in Section 2.3.1.
@@ 3.1 Lattice translation and CRT
$T\,\vec{S}_{j}\,T^{-1}=\vec{S}_{j+1}\,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{R}\,\vec{S}_{j}\,\mathcal{R}^{-1}=\vec{S}_{-j}\,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{T}\,\vec{S}_{j}\,\mathcal{T}^{-1}=-\vec{S}_{j}\,.$ (3.1)
Here $T$ and $\mathcal{R}$ are unitary operators, while the time-reversal operator $\mathcal{T}$ is anti-unitary, i.e., $\mathcal{T}i\mathcal{T}^{-1}=-i$. Explicitly, $\mathcal{T}=\left(\prod_{j=1}^{L}iY_{j}\right)\mathcal{K}$, where $\mathcal{K}$ is the complex conjugation. [^15] The anti-unitarity of $\mathcal{T}$ forces the minus sign in its action on $\vec{S}_{j}$, such that the commutation relations of $\vec{S}_{j}$ are preserved. Related to that, $\mathcal{T}$ acts on each qubit such that its square is 1-1, i.e., each qubit is a Kramers’ doublet. On the other hand, $T$ and $\mathcal{R}$ are products of the SWAP gates. These operators and their LSM constraints were discussed in [77, 78, 79, 64]. Below, we will discuss a related, but different anomaly associated with lattice translation $T$ and the following anti-unitary operator
@@
$T^{L}=1\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \Pi^{2}=(-1)^{L}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ T\Pi=\Pi T^{-1}\,.$ (3.4)
The important $(-1)^{L}$ phase follows from the fact that each qubit is a Kramers’ doublet (see footnote 15). For completeness, we also record the larger algebra generated by $\mathcal{R}$, $\mathcal{T}$, and $T$:
@@ 3.2.1 Anti-ferromagnetic phase
${\cal L}={f\over 2}\partial_{\mu}\vec{n}\cdot\partial^{\mu}\vec{n}+\theta\Gamma\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \vec{n}\cdot\vec{n}=1$ (3.11)
@@
$\mathsf{C}\,\vec{n}(t,x)\,\mathsf{C}^{-1}=-\vec{n}(t,x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathsf{R}\,\vec{n}(t,x)\,\mathsf{R}^{-1}=\vec{n}(t,-x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathsf{T}\,\vec{n}(t,x)\,\mathsf{T}^{-1}=-\vec{n}(-t,x)\,.$ (3.12)
@@
In the quantum theory, the classical field $\vec{n}$ becomes an operator, which is not constrained by $\vec{n}^{2}=1$. (In order not to clutter the equations, we will not distinguish between the classical field and its quantum counterpart.) For $\theta\in\pi(1+2{\mathbb{Z}})$, this continuum theory flows to the $su(2)_{1}$ CFT, which is the same as the IR limit of the lattice model (3.9). The lattice operators $\vec{S}_{j}$ flow to the continuum operators [14, 80]
@@ 3.2.2 Ferromagnetic phase
$\mathsf{R}\,\vec{n}(t,x)\,\mathsf{R}^{-1}=\vec{n}(t,-x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathsf{T}\,\vec{n}(t,x)\,\mathsf{T}^{-1}=-\vec{n}(-t,x)$ (3.17)
@@ 3.3.1 Anomaly for ${\tilde{\Pi}}=\mathcal{R}\mathcal{T}$
${\tilde{\Pi}}^{2}=-1\,,\quad\text{odd}\penalty 10000\ L\,.$ (3.19)
@@ 3.3.2 Anomaly between translation and $\Pi$ (the lattice CRT)
$\Pi^{2}=-1\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{odd }L\,$, (3.22)
@@ 3.4 Lattice realization of the Smith isomorphism
As discussed in Appendix F.4, this deformation gaps the bulk of the two halves of the circle. Indeed, under the identification $(-1)^{j}\vec{S}_{j}\cdot\vec{S}_{j+1}\to\Gamma$ [14, 80] (see Appendix F.2), (LABEL:HDWs) corresponds to changing the $\theta$-angle by the opposite amounts away from $\theta=\pi$ on the left and right halves of the circle.
@@ 4.1 The model and its symmetries
We are particularly interested in a $\mathbb{Z}_{2}^{\mathcal{U}}$ subgroup of the internal symmetry, which is generated by [^21] When $L=0$ mod 4, one can perform a unitary transformation $\prod_{j=1,2\penalty 10000\ \text{mod}\penalty 10000\ 4}X_{j}$ to map $\mathcal{U}$ to $\prod_{j=1}^{L}\mathrm{CZ}_{j,j+1}\prod_{j=1}^{L}X_{j}$, which is sometimes known as the CZX symmetry [95].